I'm so close to solving this question but any input would be so greatly appreciated! I'm having trouble coming up with the right 3 equations for this problem:
"The sum of three numbers is 14. The first number minus three times the third number is the second number. The second number is six more than the first number. Find the three numbers."
Thank you for any input.

Answer by Boreal(15235)

x+y+z=14
x-3z=y
y=x+6
x-3z=x+6
-3z=6
z=-2
x+y=16, going back to the first equation.
y=x+6
so x+x+6=16
2x=10
x=5
y=11
The numbers are 5, 11, -2
They add to 14
The first number minus 3 (-2)= 11
The second number is 6 more than the first.